an angle is 15 more than the other. if their average is added to both angles they would be supplementary. find the measure of the angles..
Answer by mananth(16946)

x
x+15
average = (x+x+15)/2= (2x+15)/2
(2x+15)/2+x+(2x+15)/2+(x+15)=180
(4x+30)/2 +2x+15=180
((4x+30)+4x)/2=180-15
((4x+30)+4x)/2=165
((4x+30)+4x)=330
8x+30=330
8x=300
x=37.5
37.5deg & 52.5 deg are the angles
m.ananth@hotmail.ca